OpenStudy (deadshot):

Check my work please? What is the simplified form of the quantity 12 z-squared plus 25z plus 12 over the quantity 3 z-squared minus 2z minus 8 ?

OpenStudy (deadshot):

\[\frac{ 12z^2+25z+12 }{ 3z^2-2z-8 }\]

OpenStudy (deadshot):

\[\frac{ (4z + 3)(3z + 4) }{ (3z + 4)(z - 2) }\]

OpenStudy (deadshot):

Simplified it's\[\frac{ (4z+3) }{ (z-2) }\]right?

OpenStudy (deadshot):

@ParthKohli Is this correct?

OpenStudy (deadshot):

@satellite73 Is this correct?

OpenStudy (deadshot):

@kausarsalley Is this correct?

OpenStudy (anonymous):

yh...it is absolutely correct!

OpenStudy (deadshot):

Thanks! I have to make sure because I'm practicing for the dreaded third quarter finals

OpenStudy (anonymous):

ohh okay.....you are welcome!